### v1

# TAUDiff: Improving statistical downscaling for extreme weather events using generative diffusion models

Abstract
Deterministic regression-based downscaling models for climate variables often suffer from spectral bias, which can be mitigated by generative models like diffusion models. To enable efficient and reliable simulation of extreme weather events, it is crucial to achieve rapid turnaround, dynamical consistency, and accurate spatio-temporal spectral recovery. We propose an efficient correction diffusion model, TAUDiff, that combines a deterministic spatio-temporal model for mean field downscaling with a smaller generative diffusion model for recovering the fine-scale stochastic features. We demonstrate the efficacy of this approach on downscaling atmospheric wind velocity fields obtained from coarse GCM simulations. Our approach can not only ensure quicker simulation of extreme events but also reduce overall carbon footprint due to low inference times.
1 Introduction
Weather extremes are on the rise due to accelerated climate change [1]. Given their potential to severely damage life and property, it is becoming increasingly important to estimate their frequency, associated risks and economic losses beforehand [2, 3, 4]. By insuring for such losses, we can become more resilient towards extreme events [5]. Climate risk modeling often relies on historical Earth system observations [6] or physics-based general circulation models (GCMs) [7] to generate climate projections. Typically, GCMs operate at a coarse resolution ($O(10)-O(10^{2})$km) due to compute limitations. This leads to incorrect characterization of weather extremes. In recent years, machine-learning based statistical downscaling approaches have been explored to obtain realistic well-resolved climate data over specific regions [8, 9, 10]. These methods leverage historical Earth system observation data to create a non-linear mapping from bias-corrected coarse GCM simulations to the desired higher-resolution outputs.
While deterministic regression models effectively capture large-scale features, they struggle with fine-scale stochastic atmospheric processes due to low-frequency spectral bias [11]. This limitation has recently led to the adoption of generative models like GANs [10, 12, 13], and denoising diffusion models for downscaling tasks [14, 15]. Denoising diffusion models [16, 17, 18] are particularly promising due to their stability in training, reliable convergence, and high output quality. However, sampling is often time consuming. Addressing this, Karras et al. [19] explored the design space of such diffusion models, and proposed the elucidated diffusion model (EDM) which successfully reduced the number of model evaluations (from $O(10^{3})$ to $O(10)$) required to generate a single sample. Motivated by this, a correction diffusion model (CorrDiff) [14] was proposed for kilometer-scale downscaling. CorrDiff combined a UNet-based deterministic model to map the mean field and an EDM correction to capture fine-scale stochastic content.
In the context of extreme-event simulation, it is vital that both short- and long-term event statistics of downscaled data be consistent with historical observations. So, the lack of temporal modeling in downscaling models may affect dynamical consistency of downscaled data (e.g. distorted propagation of storm fronts). One could address this issue by borrowing techniques from video generation/prediction [20], as explored by Yoon et al. [21] for regional weather forecasting. However, such techniques have not yet been explored for downscaling. Moreover, large models as in  [14] are computationally intensive to train and infer. This prohibits the generation of even relatively small ($O(10^{3})$) extreme-event datasets, which are crucial for accurately quantifying climate tail risk. Given a good mean-field model, it is possible that a smaller and computationally efficient diffusion model would suffice. This would reduce overall computational demands, inference times, and improve efficiency for real-time use.
To address these challenges, we propose a computationally efficient Temporal Attention Unit enhanced Diffusion model (TAUDiff) that integrates (a) a video prediction model for dynamically consistent mean-field downscaling, and (b) a smaller guided denoising diffusion model for stochastically generating the fine-scale features. We train the models on atmospheric wind fields obtained from reanalysis dataset. The performance of TAUDiff is first compared against separate mean-field regression and end-to-end diffusion models under a fixed training budget. Finally, we evaluate the downscaling performance of TAUDiff on an ensemble of bias-corrected coarse GCM outputs using various spectral statistics. We finally discuss its potential in producing accurate and computationally efficient extreme-event datasets, and the implications of model inference times and carbon footprint offsetting.
2 Methods
Overview
We demonstrate the efficacy of TAUDiff in downscaling atmospheric wind fields over the European region. For our training, we use the atmospheric reanalysis dataset (ERA5) at $0.25^{\circ}$ lat-lon resolution produced by the European Center for Medium-range Weather Forecasts (ECMWF) [6]. Instead of a single time instance input like in [14], our approach uses a deterministic regression component that takes a temporal sequence of coarsened ERA5 wind velocity snapshots with orography data as input. Here, the high-resolution ERA5 wind fields from the final time step of the sequence serves as the target. Instead of using coarse interpolation, we use lowpass spherical wavelet filtering [22, 23] to create band-limited low-resolution ERA5 fields to ensure proper scale separation. This approach closely mirrors real-world scenarios where bias-corrected GCM data lacks fine-scale spatio-temporal features. We use the Community Atmosphere Model 4.0 (CAM4) [24] (at $1^{\circ}$ lat-lon resolution) as the coarse GCM in this study. We briefly discuss the mean-field and diffusion model components of TAUDiff below.
Figure 1: Schematic of the TAUDiff model
TAUDiff Component 1: Mean model
We adopt the Simple yet better Video Prediction (SimVP) architecture [25] consisting of a spatial backbone, and a translator for temporal modelling, ensuring temporal coherence and simplicity as compared to the more complex transformer based architectures [20]. We specifically consider a UNet for the spatial backbone, and the temporal attention unit (TAU) [26] for the translator. The TAU first independently models spatial dependency via static, and both cross-channel and temporal dependencies using dynamical attention units, respectively, and then combines them. We train the mean model using a weighted combination of mean absolute error (MAE), mean squared error (MSE), and to additionally maintain dynamical consistency, physics-based losses on advection ($\boldsymbol{u}\cdot\nabla\boldsymbol{u}$), vorticity (×$\nabla\times\boldsymbol{u}$) and divergence ($\nabla\cdot\boldsymbol{u}$) of wind fields ($\boldsymbol{u}$) are also considered. Although dynamically consistent predictions are possible with this mean model, the downscaled fields still lack the stochastic fine scale features. This is where Component 2 comes into play.
TAUDiff Component 2: Correction diffusion model
To capture the residual stochastic fine scale features (which cannot be captured by the mean model), we build a relatively small correction diffusion model ($\sim O(1)$ million (M) parameters) trained using a score-matching loss [27]. To maintain consistency of our approach, we use a SimVP architecture as in the mean model but with a residual dense UNet as the spatial backbone. Once the model is trained, a data sample can be generated by solving a stochastic differential equation modelling a reverse diffusion process [17, 19]. Since the conditional input to the diffusion model is the mean model output (for a single time instance), the TAU morphs into a Channel Attention Unit (CAU) where the dynamical attention unit now models cross-channel dependencies and their relative importance (see figure 1 for a detailed schematic of TAUDiff.).
3 Results
Experimental protocol
We demonstrated the potential of our framework by comparing three models: a deterministic mean-field regression, an end-to-end diffusion, and our TAUDiff model, each with $O(10)$M trainable parameters overall. These models were trained over 40 years of ERA5 atmospheric wind data over Europe (1980-2020) (figure 2(a)), and validated over 2021-23. All the models were trained on a single T4 GPU over 50 epochs, with training times of 24, 48, and 60 hours for the mean, end-to-end diffusion, and the TAUDiff models, respectively.
Figure 2: (a) European region used for training the downscaling models, with select locations used for evaluating performance. Comparison of model predictions: (b) vorticity snapshot at UTC: 2023-12-31 21:00, (c) spatial spectrum, and (d) temporal spectra at select locations shown in (a).
Validation
Qualitatively, the vorticity contour predictions of mean and TAUDiff models demonstrate dynamical consistency of storm fronts, whereas the end-to-end diffusion model distorts them due to noise injection (see the circled zone in figure 2(b)). Quantitatively, pointwise statistics computed over validation years 2021-23 show good recovery of spatial and temporal spectrum for TAUDiff, while mean model underrepresents, and end-to-end diffusion overrepresents higher temporal frequencies, respectively (see figure 2(c) and 2(d)).
Testing
The performance of our TAUDiff model was then evaluated on downscaling bias corrected CAM4 obtained wind fields over 40 years. Bias correction is done by quantile-mapping [28] the 40-year distribution of each grid cell to that of ERA5, wavelet-filtered to GCM resolution. As earlier, we obtain physically consistent output, and remarkable spectral recovery (see figures 3(a-c)). Although only a simple quantile mapping [28] is adopted for bias correction, we see good agreement with ERA5 ground truth in the local storm counts (see figure 3(d)). This cements the need for a stochastic correction using a diffusion model for accurate extreme-event risk estimation.
Figure 3: Assessment of downscaling performance on bias corrected CAM4 data: (a) Vorticity contours at a representative time instance, (b) temporal spectrum, (c) vorticity distributions, and (d) local storm counts.
4 Discussion
Overall, our proposed video-prediction-based TAUDiff model demonstrates dynamically consistent downscaling, remarkable reconstruction of spatio-temporal fine scale features, and viable inference times with the use of a small correction diffusion model. Since coarse and fine scale content of the atmospheric fields are resolved well, accurate estimation of storm statistics was possible. Thus, we show that TAUDiff has immense potential in accurate, and computationally efficient estimation of extreme weather events. Accurate estimation of extreme events would enable reliable risk estimations and realistic climate projections. Our future work would involve staging of TAUDiff models to obtain multi-resolution outputs for extreme weather event simulations while maintaining reasonable inference times.
Smaller models with low inference times also mean a lower carbon footprint. In our TAUDiff architecture with the size of $O(10)$M parameters, the diffusion model component is only $O(1)$M parameters. This allows for efficient inference and enables operationalization at scale. For inference on just one year worth of 3-hourly resolved wind fields over Europe, it takes approximately 30 minutes on a single T4, and about 4 minutes on an H100 GPU. In contrast, the end-to-end diffusion model ($O(10)$M parameters) takes about 80 minutes on a T4 GPU, and approximately 9 minutes on an H100 GPU. In large-scale operational settings, such as querying TAUDiff millions of time to create large exteme-event datasets, strategies for offsetting the carbon footprint should be considered. One option is to run inference on different cloud locations; for instance, 100 hours on an A100 GPU based AWS EC2 instance located in Canada (Central) can produce 0.5kg $\text{CO}_{2}$q., fully offset by renewable energy; whereas, the same located in the US (North Virginia) can produce 9.25kg $\text{CO}_{2}$q. with no offset at all (estimations made using the Machine Learning Impact calculator [29]).